# Single-View Tomographic Reconstruction Using Learned Primal Dual

### v1

@@ I Introduction
where $\mathcal{R}$ is the regularization penalty functional. Examples include Tikhinov [daun2005solution], Total Variation (TVmin) [asaki2006abel], higher-order TVmin [chan2015high], and $L1/L2$ of the gradient [breckling2023box]. Such techniques are typically approached by solving their respective loss models using a direct numerical optimization technique, e.g. the Primal Dual approach of Chambolle & Pock [chambolle2011first]. This approach has been studied extensively in literature, even for single-view tomographic reconstruction [zhang2018stability].
@@
Herein we consider a Learned Primal Dual (LPD) architecture for single-view tomographic reconstruction problems. This method has recently shown promise in limited-angle and extreme few-view modalities for tomographic reconstruction, particularly compared to direct numerical optimization-based methods in the severely under-determined cases [Adler2019]. The LPD scheme is a so-called unrolled scheme, as it is constructed by substituting the eponymous Primal and Dual steps with shallow CNNs. These shallow networks are coupled in sequence, with the forward model $\mathcal{A}$ and adjoint $\mathcal{A}^{*}$ applied between the primal and dual steps, iteratively enforcing the physical acquisition model constraints. This architecture significantly reduces the network complexity compared to monolithic methods, and ensures reconstructions remain consistent with the projection data, minimizing the hallucination risks common to the post-processing approaches. These are key safeguards for our application, given the severe acquisition model constraints in axisymmetric tomography.
The use of NNs as part, or all, of the inversion procedure in single-view modalities is not new. Multilayer Perception (MLP) models were implemented with success as early as 1992 [bishop1992neural], with NN-based methods arriving in the literature by 2002 [Ma2002]. However, the rapid growth of available software tools, computational resources, and broad interest in machine learning (ML) has resulted in a significant uptick in new studies [RODRIGUEZ2021119011, Duann2023, liu2023physics].
@@ II-A Training Data
Given a particular surrogate imaging apparatus model $\mathcal{F}:X\rightarrow Y$, each scene in $X$ is constructed by combining one or more randomly-selected axisymmetric target objects, and assigning a volumetric density distribution. The corresponding projection image in $Y$ is then generated by calling the forward model $\mathcal{F},$ along with any associated distortion model (e.g. blur, and noise) $K$. Forward models denoted by $\mathcal{F}$ are either an analytic, closed-form model akin to those denoted below in Table I, or an absorption image generated using a physics ray-tracing engine like gVirtualXray (gVXR) [vidal2024x], Monte Carlo N-Particle (MCNP) [mcnp], or HADES [aufderheide2010inclusion]. We note that models $\mathcal{F}$ are only used to generate training data, and are distinct from the discretized models discussed in the following subsection.
@@ II-B Discretized Forward Models
The usual approach to discretizing single-view, axisymmetric projection operators is not dissimilar from those taken in the more-general Radon transform case [averbuch2008framework], apart from the reduction from a 3D target space the a 2-dimensional cylindrical space reflected in (1). We consider the complete radiographic projection in both modalities. This choice is a departure from typical discretizations of (1), given that the idealized reconstruction problem is fully determined with precisely half of a complete projection of the cylindrical scene. In practice, the additional samples in the over-determined problem can help suppress the effects of measurement noise in the reconstruction process, but are often handled post-hoc.
@@
where $n>0$ is the number of rows present in the ROI, and $\otimes$ is the Kroneker product.
@@ II-C Network Architecture
All models were implemented and trained using the PyTorch deep learning framework [paszke2019pytorch]. Each convolutional layer within the Primal and Dual CNN blocks, as depicted in Figure 2, utilized a 3x3 kernel followed by a Parametric Rectified Linear Unit (PReLU) activation function. The final layer of each block uses a linear activation. PREeLU if often selected over standard ReLU to allow the network to learn and adapt the negative slope; as was the case in [Adler2019]. In the usual way, these networks are trained end-to-end using Adam against an MSE loss, and a fixed learning rate of $10^{-4}.$
@@ II-D Evaluation Metrics
In the sections that follow, we utilize several common benchmarking metrics– namely, the Mean-Square Error (MSE), Normalized Mean-Square Error (NMSE), the Structural Similarity Index Metric (SSIM), and Pseudo Signal to Noise Ratio (PSNR). Given $N$, the total number of pixels in the image space $X$, a known non-zero ground truth $u\in X$ and approximation $v\in X$, we write
@@ III-B Cone-beam Modality
Similarly, a model for the total blur of the system was estimated using a Tungsten rollbar, fitting a parameterized regressor to the edge-spread function (ESF) following [schach2002using]. The resulting point-spread function (PSF) kernal $K$ is assumed to be radially symmetric, fully determined by the parameters ${\bf a}=(a_{1},\ldots,a_{6})$ such that
@@ III-D Results
Mirroring the results found in the parallel-beam case, the LPD solver outperformed the included classical solvers; with similar caveats. We selected FDK [feldkamp1984practical] as the unregualrized baseline reconstruction example. This is a fast method that is similar to Gridrec, but can directly accommodate the cone-beam acquisition. As before, we include results from both the Tikhonov and TV-min methods, again identifying their respective regularization parameters $\lambda$ using a grid search of the evaluation data.

### v2

@@ I Introduction
where $\mathcal{R}$ is the regularization penalty functional. Examples include Tikhonov [daun2005solution], Total Variation (TVmin) [asaki2006abel], higher-order TVmin [chan2015high], and $L1/L2$ of the gradient [breckling2023box]. Such techniques are typically approached by solving their respective loss models using a direct numerical optimization technique, e.g. the Primal Dual approach of Chambolle & Pock [chambolle2011first]. This approach has been studied extensively in literature, even for single-view tomographic reconstruction [zhang2018stability].
@@
Herein we consider a Learned Primal Dual (LPD) architecture for single-view tomographic reconstruction problems. This method has recently shown promise in limited-angle and extreme few-view modalities for tomographic reconstruction, particularly compared to direct numerical optimization-based methods in the severely under-determined cases [Adler2019]. The LPD scheme is a so-called unrolled scheme, as it is constructed by substituting the eponymous Primal and Dual steps with shallow CNNs. These shallow networks are coupled in sequence, with the forward model $\mathcal{A}$ and adjoint $\mathcal{A}^{*}$ applied between the primal and dual steps, iteratively enforcing the physical acquisition model constraints. This architecture significantly reduces the network complexity compared to monolithic methods, and ensures reconstructions remain consistent with the projection data, minimizing the hallucination risks common to the post-processing approaches. These are key safeguards for axisymmetric tomography, given the severity of the acquisition model assumptions.
The use of NNs as part, or all, of the inversion procedure in single-view modalities is not new. Multilayer Perceptron (MLP) models were implemented with success as early as 1992 [bishop1992neural], with NN-based methods arriving in the literature by 2002 [Ma2002]. However, the rapid growth of available software tools, computational resources, and broad interest in machine learning (ML) has resulted in a significant uptick in new studies [RODRIGUEZ2021119011, Duann2023, liu2023physics].
@@ II-A Training Data
Given a particular surrogate imaging apparatus model $\mathcal{F}:X\rightarrow Y$, each scene in $X$ is constructed by combining one or more randomly-selected axisymmetric target objects, and assigning a volumetric density distribution. The corresponding projection image in $Y$ is then generated by calling the forward model $\mathcal{F},$ along with any associated distortion model (e.g. blur, and noise) $K$. Forward models denoted by $\mathcal{F}$ are either an analytic, closed-form model akin to those denoted below in Table I, or an absorption image generated using a physics ray-tracing engine like gVirtualXray (gVXR) [vidal2024x, vidal2025x], Monte Carlo N-Particle (MCNP) [mcnp], or HADES [aufderheide2010inclusion]. We note that models $\mathcal{F}$ are only used to generate training data, and are distinct from the discretized models discussed in the following subsection.
@@ II-B Discretized Forward Models
The usual approach to discretizing single-view, axisymmetric projection operators is not dissimilar from those taken in the more-general Radon transform case [averbuch2008framework], apart from the reduction from a 3D target space to a 2-dimensional cylindrical space reflected in (1). We consider the complete radiographic projection in both modalities. This choice is a departure from typical discretizations of (1), given that the idealized reconstruction problem is fully determined with precisely half of a complete projection of the cylindrical scene. In practice, the additional samples in the over-determined problem can help suppress the effects of measurement noise in the reconstruction process, but are often handled post-hoc.
@@
where $n>0$ is the number of rows present in the ROI, and $\otimes$ is the Kronecker product.
@@ II-C Network Architecture
All models were implemented and trained using the PyTorch deep learning framework [paszke2019pytorch]. Each convolutional layer within the Primal and Dual CNN blocks, as depicted in Figure 2, utilized a 3x3 kernel followed by a Parametric Rectified Linear Unit (PReLU) activation function [he2015delving]. The final layer of each block uses a linear activation. PREeLU is often selected over standard ReLU to allow the network to learn and adapt the negative slope; as was the case in [Adler2019]. In the usual way, these networks are trained end-to-end using Adam [adam2014method] against an MSE loss, and a fixed learning rate of $10^{-4}.$
@@ II-D Evaluation Metrics
In the sections that follow, we utilize several common benchmarking metrics– namely, the Mean-Square Error (MSE), Normalized Mean-Square Error (NMSE), the Structural Similarity Index Metric (SSIM) [wang2004image], and Pseudo Signal to Noise Ratio (PSNR). Given $N$, the total number of pixels in the image space $X$, a known non-zero ground truth $u\in X$ and approximation $v\in X$, we write
@@ III-B Cone-beam Modality
Similarly, a model for the total blur of the system was estimated using a Tungsten rollbar, fitting a parameterized regressor to the edge-spread function (ESF) following [schach2002using]. The resulting point-spread function (PSF) kernel $K$ is assumed to be radially symmetric, fully determined by the parameters ${\bf a}=(a_{1},\ldots,a_{6})$ such that
@@ III-D Results
Mirroring the results found in the parallel-beam case, the LPD solver outperformed the included classical solvers; with similar caveats. We selected FDK [feldkamp1984practical] as the unregularized baseline reconstruction example. This is a fast method that is similar to Gridrec, but can directly accommodate the cone-beam acquisition. As before, we include results from both the Tikhonov and TV-min methods, again identifying their respective regularization parameters $\lambda$ using a grid search of the evaluation data.